please help me with this equation
3t+1/2t=28 solve and check
can you explain how you get the answer



Answer by Socom491(19)

Ok, so first we have to combine the like terms, which are 3t and 1/2t, when we do this we get 3 and 1/2t, so now we plug it in to the equation.
3 1/2t=28
/3 1/2= / 3 1/2
And...28 divided by 3.5 equals 8. So t=8.
To check we plug eight in replacement of t's.
3(8)+1/2(8)=28...and it checks.